OpenStudy (anonymous):

For the Partial Fraction , Parts (Denominator - Repeated Linear Factor ) , how to prove the biggest power of x^3 in the denominator ? For example P(x)+Q/(x-3)^3 , how to prove to A/x-3 + B/(x-3)^2 + C/(x-3)^3

OpenStudy (shamil98):

P(x)+Q/(x-3)^3 = A/x-3 + B/(x-3)^2 + C/(x-3)^3 ?

OpenStudy (anonymous):

yaya

OpenStudy (shamil98):

are you given any more information?

OpenStudy (anonymous):

nope just wanna show that P(x)+Q/(x-3)^3 = A/x-3 + B/(x-3)^2 + C/(x-3)^3

OpenStudy (shamil98):

x >= 4 so we don't get undefined through division i guess. so if x = 4 P(4) + Q = A + B + C

OpenStudy (anonymous):

u wait i am going to scan somethg

OpenStudy (anonymous):

..